OpenStudy (anonymous):

Can anyone please explain the topic calendars ? eg: What was the day on 15th august 1947 ? Explain step by step - Intra month , inter month , inter year. Thanks.

OpenStudy (anonymous):

like.......17th June, 1998 = (1997 years + Period from 1.1.1998 to 17.6.1998) Odd days in 1600 years = 0 Odd days in 300 years = (5 x 3) 1 97 years has 24 leap years + 73 ordinary years. Number of odd days in 97 years ( 24 x 2 + 73) = 121 = 2 odd days. Jan. Feb. March April May June (31 + 28 + 31 + 30 + 31 + 17) = 168 days 168 days = 24 weeks = 0 odd day. Total number of odd days = (0 + 1 + 2 + 0) = 3. in this way u can find for ur date

OpenStudy (anonymous):

Thanks but I dint get anything :p But some other short cut ( easier way to understand) ?

OpenStudy (anonymous):

oww wait let me think........

OpenStudy (anonymous):

If you could find some video related to it , it would be great !

OpenStudy (experimentx):

the Gregorian calendar has same number of days in every month ... except for February (29) total days = difference of today to 15 th august + 365.4*(2013 - 1947) 15 th august 1947 day = today's day + (total days) mod 7

OpenStudy (anonymous):

Is it in accordance with the standard english calendar @experimentX ?

OpenStudy (experimentx):

yes it is .. check every calendar, all months have the same number of days except February

OpenStudy (anonymous):

My teacher told me a different method to calculate. He assumed that if all the months had 28 days then the extra days in each month would be : jan 3 , feb 0/1 , march 3 , april 2 , may 3 and so on ! Then he used this data to calculate the day on a particular date ( inter month ) . Iam not able to understand this method , if you get to know , please let me know @experimentX !

OpenStudy (experimentx):

it is essentially same ... if all months were 28 then it's natural that extra days would be jan 3 , feb 0/1 , march 3 , april 2 , may 3 ... and weeks repeat after 7 days, if it's dec last and it's sunday then ... feb 1 would be => 3+1 = thursday .. same principle ... it shortens the number.

OpenStudy (anonymous):

that is alrite. But if for solving , in entrance papers , I need to solve via shortcuts. Please help in inter year section as well.

OpenStudy (anonymous):

* but for solving

OpenStudy (experimentx):

find the extra days in each year (total)

OpenStudy (anonymous):

lets do an example. its given that july 13 2012 is a friday , what day will it be on 17'th december 2012 ?

OpenStudy (anonymous):

please explain step by step , by the method I mentioned above

OpenStudy (experimentx):

no ... assume everymonth as 28 days, find the extra days in every month

OpenStudy (anonymous):

no ?

OpenStudy (anonymous):

What after finding extra days ?

OpenStudy (experimentx):

well ... add 0.25 to feb

OpenStudy (anonymous):

It would have been gr8 , if you could explain the basic method involved step by step.

OpenStudy (experimentx):

okay .. give me the data first.

OpenStudy (anonymous):

why 0.25 ?

OpenStudy (anonymous):

its given that july 13 2012 is a friday , what day will it be on 17'th december 2012 ?

OpenStudy (experimentx):

because of leap year.

OpenStudy (anonymous):

see what I have followed from my class is that , I will start adding extra days from july till november ( not of december because its effect will be felt on jan ) . Feb is out of question because my data varies from july to december. Had it been jan , then concept of leap year would have taken into account. Now after adding total number of days = 13 , the calender will shift by 13 days. It means that we'll calculate the day on 13+13=26 july , ( adding 7 days 4times , will yield a friday again then -2days ) which will be a Wednesday. Now 26 december will be Wednesday , so 17th dec will come out to be saturday ! But its wrong ! It will be monday! Please help !

OpenStudy (experimentx):

well ... i'm not sure about that. can you count from March to Jan?

OpenStudy (anonymous):

that would be inter year ! Here iam discussing inter month. Both are slightly diff frm each other

OpenStudy (experimentx):

not sure how you would do.

OpenStudy (anonymous):

:-( Anyways thanks for your dedicated efforts ! :-)